# Physics of droplet regulation in biological cells

### v1

@@ 4.2 Wetting of filaments
Under full wetting conditions, proteins may condense around a filament to form a cylindrical layer of droplet material with uniform thickness. Over time, however, surface tension may cause this layer to break up into a series of smaller beaded droplets with finite spacing, via the Plateau–Rayleigh instability (see figure 4.2A). Recent works have demonstrated this effect for certain microtubule-binding proteins [334, 335]. The spacing of beaded droplets increases with the original thickness of the wetted cylinder [334] and then in turn controls the spacing of subsequent microtubules which nucleate from these droplets.
If multiple filaments are simultaneously wetted by the same droplet, the capillary bridges that form between the filaments give rise to adhesive forces. These adhesive forces then lead to bundling [336] and repositioning [337] of adjacent filaments. Adhesive forces between droplets and filaments can also control the positioning of droplets within the cell, concentrating them within filament-rich regions [260, 338]. Droplet position and shape may also be regulated by active filament polymerisation within the droplet, resulting in droplet translocation and deformation [5]. Finally, droplet formation on individual strands of semi-flexible filaments can result in a co-condensation process, in which the filament is pulled into the condensate by capillary forces. This effect can be resisted if the filament is held at a sufficiently high tension which opposes the capillary pull [339], and provides an additional method of force generation within the cell.
@@ 4.3 Interaction with membranes
Cellular membranes take the form of phospholipid bilayers, and allow the cell to partition different regions of space. Membranes are typically very thin and flexible, and contain an array of proteins and other biomolecules which reorganise laterally in a fluid-like manner. Droplets come into contact with many different membranes within the cell, including the plasma membrane, membrane-bound organelles, vesicles, and the endoplasmic reticulum. Droplet–membrane interactions are mediated by both physical (e.g., wetting) and biomolecular (e.g., binding) effects [340, 341, 342, 60]. In this section, we discuss interactions between macroscopic droplets and membranes, as well as the possibility of phase separation phenomena within the membrane itself.
@@ 4.3.2  Condensation within membranes
Despite extensive work investigating the membrane-bound phase separation of lipid domains in model membranes, it remains unclear whether such phenomena occur in intact biological cells. Recent work, however, suggests that membrane heterogeneities result from correlations induced by the proximity of the membrane to a critical point [355]. As such, the ability of a lipid bilayer to phase separate affects membrane organisation, regardless of the fact that macroscopic phase separation may not occur. A detailed discussion of the role of critical phenomena in cellular membranes is provided by [355].
@@ 4.4 Droplets in elastic meshes
The interior of the cell is criss-crossed by a series of elastic meshes, such as the cytoskeleton in the cytoplasm, and chromatin in the nucleus. In order to grow larger than the characteristic pore size of these meshes, droplets must necessarily interact with the mesh, which may deform as a result. Droplet–mesh interactions can lead to restricted droplet growth [361], suppressed droplet coalescence [361, 362], sub-diffusive droplet motion [363], and spatial localisation of droplets [221]. The presence of meshes may also prevent condensates from settling in very large cells in which gravity matters [78]. In turn, the formation of droplets may affect the mechanical properties of the mesh itself [364, 365].
@@ 4.4.1  Network-excluding droplets
To capture the dynamics of droplets embedded in elastic meshes, we describe the dilute and dense phases separately, as in section 2.2. If there exist multiple droplets, then the evolution of droplet sizes is given by equation (2.44), but with the equilibrium volume fraction of the dilute phase, $\phi_{\textrm{out}}^{\textrm{eq}}$, now a function of $P_{\textrm{el}}$, which in turn depends on the size of the droplet. In general, $P_{\textrm{el}}$ can also depend on space, as discussed in section 4.4.3 below. Note that the presence of the mesh restricts droplet motion, such that droplets can be assumed to be fixed in place once they have nucleated. The evolution of the dilute phase follows equation (2.43).
@@ 4.4.2  Network permeation
If the elastic network is sufficiently stiff and the mesh size is sufficiently large, it may be energetically favourable for droplets to permeate, rather than exclude, the network. As droplets permeate the network, a further contribution to the free energy is induced, resulting from the differential surface energies of network–droplet and network–solvent interfaces. Following the work of Ronceray et al. [366], we capture this contribution by introducing a ‘wetting energy’ of the form
@@ 4.4.3  Complex properties of the mesh
Equation (4.10) highlights the dependence of the equilibrium droplet size on the elastic properties of the mesh, via the elastic pressure. Typically, droplets will be larger in softer materials. If the elastic mesh has long-range variations in its stiffness, we would thus expect larger droplets to form in softer regions of the mesh [134, 378]. Over long timescales, it is then energetically favourable for droplets in stiffer regions to dissolve in favour of droplet growth in softer regions, in order to equilibrate chemical potential across the system. This process has been termed elastic ripening [143], and is distinct from the classical Ostwald ripening effect. In particular, it has been shown that larger droplets can dissolve at the expense of smaller droplets for appropriate stiffness gradients, in contrast to Ostwald ripening [143].
@@ 5 Chemically active droplets
Cells control the properties of their biomolecules by chemical modifications, such as post-translational modifications in the case of proteins. Such modifications can affect physical interactions with other biomolecules and thus phase separation. Many condensates in cells are regulated by these modifications [382, 220, 53, 219, 43], including centrosomes [383] and stress granules [384]. To develop a physical understanding of these regulatory mechanisms, we conceptualise biomolecular modifications as chemical transitions between two forms of the same molecule, differing in their interaction parameters encoded by $w_{i}$ and $\chi_{ij}$. After introducing a general framework for describing (active) chemical reactions, we couple reactions to phase separation, identifying two distinct classes of chemically active droplets in the simplest case of a binary mixture.
@@ 5.1 Chemical reactions of many components
Reaction kinetics describe how the mixture composition $\Phi=(\phi_{1},\ldots,\phi_{N_{c}})$ changes over time. For a homogeneous mixture, there are no diffusive fluxes, implying $\partial_{t}\phi_{i}=s_{i}$, where $s_{i}=\sum_{\alpha=1}^{N_{\mathrm{r}}}\sigma^{(\alpha)}_{i}s^{(\alpha)}$ is the net production of component $i$. The net flux $s^{(\alpha)}$ of reaction $\alpha$ can be split in a forward and a backward direction, $s^{(\alpha)}=s^{(\alpha)}_{\rightarrow}-s^{(\alpha)}_{\leftarrow}$, which must obey detailed balance of the rates [23, 385],
@@
Detailed balance of the rates ensures that the reaction ceases ($s^{(\alpha)}=0$) when chemical equilibrium as described by equation (5.2) is reached, and also that the reaction proceeds in the forward direction when products are energetically favoured over reactants. A simple theory that obeys these constraints is transition state theory (see appendix F and ref. [386, 387, 388, 389, 390, 87]),
$s^{(\alpha)}=k_{\alpha}\sinh\biggl{[}-\frac{\sum_{i}\sigma^{(\alpha)}_{i}\,\mu_{i}}{2k_{\mathrm{B}}T}\biggr{]}\;,$ (5.4)
@@
where the square bracket describes energetic constraints, whereas kinetics are captured by the positive rate coefficient $k_{\alpha}$, which can depend on composition.
@@
The dynamics described by equation (5.4) generally drive the system towards equilibrium. However, biological systems are active and driven away from equilibrium by their metabolism [391]. A theoretical representation of these metabolic processes involves the conversion of high-energy fuel molecules into low-energy waste molecules via chemical reactions. These reactions allow ordered structures within the cell to be maintained while increasing the total entropy of the cell and its environment. Note that although the energy difference between fuel and waste molecules is dissipated as heat, resulting temperature changes are typically negligible in cellular systems [98]. If fuel molecules are plentiful, they essentially provide an infinite particle reservoir with fixed chemical potential, also known as a chemostat. If the production of waste molecules does not significantly change their concentration in the environment, they can also be considered to be chemostatted. The energy difference between chemostatted species drives chemical reactions out of equilibrium [392, 393, 394]. This framework can also be used to describe replenishing systems, where waste molecules (e.g., ADP) are turned back into fuel molecules (ATP) by some mechanism (ATP synthase) whose energy input is not described. In all these cases, the supplied energy is used locally to drive processes against their thermodynamic tendency, implying that the resulting system is active matter [395].
Figure 5.1: Qualitatively different classes of chemically active droplets. Reaction flux $s$ as a function of the fraction $\phi$ of droplet material for various qualitatively different functions $s(\phi)$. The flux diverges for extreme fractions (see equation (5.5)) and has either one root $\phi_{0}$ (panel A) or three roots (panel B). Homogeneous states with fraction $\phi_{0}$ are stable with respect to reactions if $s^{\prime}(\phi_{0})<0$, implying that only the outer two roots are stable in panel B. The relative positions of these roots in comparison to the equilibrium fractions $\phi^{(0)}_{\mathrm{in/out}}$ delimiting the blue binodal region determines the basal reaction rates $\Gamma_{\mathrm{in/out}}$. This positioning leads to the two distinct classes of externally maintained droplets (orange, $\Gamma_{\mathrm{in}}<0$, $\Gamma_{\mathrm{out}}>0$) and internally maintained droplets (green, $\Gamma_{\mathrm{in}}>0$, $\Gamma_{\mathrm{out}}<0$); all other cases lead to homogeneous systems (grey lines).
@@
Chemical reactions also take place in heterogeneous systems, where the composition $\Phi$ varies in space and time. We discussed the dynamics of such systems in section 2, and the central equation (2.14) already contains the source term $s_{i}$, which accounts for chemical reactions as described by equation (5.4). The combination of these equations implies that chemical reactions affect compositions only locally, whereas spatial transport takes place via diffusion and advection.
@@ 5.2 Chemically active binary mixtures
To discuss the influence of chemical reactions on droplets, we now focus on the simple case of a binary mixture of droplet material $D$ that segregates from precursor material $P$. The system’s state is described by the volume fraction $\phi(\boldsymbol{r},t)$ of the droplet material, while precursor occupies the fraction ${1-\phi}$. Thermodynamic properties of the mixture are captured by the exchange chemical potential $\bar{\mu}$; see section 2.1.3. If the two species can freely convert into each other, ${P\rightleftharpoons D}$, the system can in principle attain the minimum free energy (where $\bar{\mu}=0$, see figure 2.1a) at every point in space, implying a homogeneous mixture as the only possible equilibrium state.
To investigate more interesting dynamics, we can drive the system out of equilibrium by introducing an additional chemical reaction, $P+F\rightleftharpoons D+W$, where the fuel $F$ and waste $W$ are chemostatted. For simplicity, we assume that fuel and waste distribute fast and are dilute, so they do not affect phase separation [396]. These components thus simply provide a chemical potential difference $\Delta\!\mu=\mu_{F}-\mu_{W}$, which can drive the active reaction everywhere. The combined reaction flux $s$ then reads,
@@
where $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are the respective reaction coefficients for the passive and active reactions, which are both positive and may depend on the composition $\phi$.
The chemical reaction flux given by equation (5.5) creates droplet material when the droplet fraction $\phi$ is very small, i.e., $s\rightarrow\infty$ in the limit of $\phi\rightarrow 0$. This occurs because the exchange chemical potential diverges ($\bar{\mu}\rightarrow-\infty$) as seen in equation (2.17) and figure 2.1b. Conversely, we find $s\rightarrow-\infty$ for $\phi\rightarrow 1$, implying that droplet material is converted to precursor if its concentration is too high. As a result, the reaction flux $s$ can be approximated as an explicit function of the composition $\phi$. The limits of the function $s(\phi)$ are constrained by thermodynamics (see figure 5.1), whereas the behaviour at intermediate volume fractions depends on the precise choice of all parameters and particularly the composition-dependence of $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$. Figure 5.1 shows various different functional forms of $s(\phi)$, which we categorise below based on the qualitative behaviour they induce. Recall from equation (2.23) that the dynamics of the binary mixture are then governed by
@@
where $\Lambda$ denotes the diffusive mobility.

### v2

@@ 4.2 Wetting of filaments
Under full wetting conditions, proteins may condense around a filament to form a cylindrical layer of droplet material with uniform thickness. Over time, however, surface tension may cause this layer to break up into a series of smaller beaded droplets with finite spacing, via the Plateau–Rayleigh instability (see figure 4.2A). Recent works have demonstrated this effect for certain microtubule-binding proteins [341, 342]. The spacing of beaded droplets increases with the original thickness of the wetted cylinder [341], and in turn controls the spacing of microtubules that subsequently nucleate from these droplets.
If multiple filaments are simultaneously wetted by the same droplet, the capillary bridges that form between the filaments give rise to adhesive forces. These adhesive forces then lead to bundling [343, 344] and repositioning [345] of adjacent filaments. Adhesive forces between droplets and filaments can also control the positioning of droplets within the cell, concentrating them within filament-rich regions [264, 346]. Droplet position and shape may also be regulated by active filament polymerisation within the droplet, resulting in droplet translocation and deformation [5]. Finally, droplet formation on individual strands of semi-flexible filaments can result in a co-condensation process, in which the filament is pulled into the condensate by capillary forces. This effect can be resisted if the filament is held at a sufficiently high tension which opposes the capillary pull [347], and provides an additional method of force generation within the cell. Droplet–filament co-condensation has also been implicated in the spatial organisation of chromatin [348], with protein–DNA co-condensates preferentially nucleating at DNA loops.
@@ 4.3 Interaction with membranes
Cellular membranes take the form of phospholipid bilayers, and allow the cell to partition different regions of space. Membranes are typically very thin and flexible, and contain an array of proteins and other biomolecules that can reorganise laterally in a fluid-like manner. Droplets come into contact with many different membranes within the cell, including the plasma membrane, membrane-bound organelles, vesicles, and the endoplasmic reticulum. Droplet–membrane interactions are mediated by both physical (e.g., wetting) and biomolecular (e.g., binding) effects [349, 350, 351, 60]. In this section, we first discuss interactions between macroscopic droplets and membranes, before considering the possibility of phase separation phenomena within the membrane itself.
@@ 4.3.2  Condensation within membranes
Despite extensive work investigating the membrane-bound phase separation of lipid domains in model membranes, it remains unclear whether such phenomena occur in intact biological cells. Recent work, however, suggests that membrane heterogeneities result from correlations induced by the proximity of the membrane to a critical point [364]. As such, the ability of a lipid bilayer to phase separate can affect membrane organisation, regardless of the fact that macroscopic phase separation may not occur. A detailed discussion of the role of critical phenomena in cellular membranes is provided by [364].
@@ 4.4 Droplets in elastic meshes
The interior of the cell is criss-crossed by a series of elastic meshes, such as the cytoskeleton in the cytoplasm, and chromatin in the nucleus. In order to grow larger than the characteristic pore size of these meshes, droplets must necessarily interact with the mesh, which may deform as a result. Droplet–mesh interactions can lead to restricted droplet growth [370], suppressed droplet coalescence [370, 371], sub-diffusive droplet motion [156], and spatial localisation of droplets [225]. The presence of meshes may also prevent condensates from settling in very large cells in which gravitational effects are relevant [78]. In turn, the formation of droplets may affect the mechanical properties of the mesh itself [372, 373].
@@ 4.4.1  Network-excluding droplets
To capture the dynamics of droplets embedded in elastic meshes, we describe the dilute and dense phases separately, as in section 2.2. If multiple droplets exist, then the evolution of droplet sizes is given by equation (2.44), but with the equilibrium volume fraction of the dilute phase, $\phi_{\textrm{out}}^{\textrm{eq}}$, now a function of $P_{\textrm{el}}$, which in turn depends on the size of the droplet. In general, $P_{\textrm{el}}$ can also depend on space, as discussed in section 4.4.3 below. Note that the presence of the mesh restricts droplet motion, such that droplets can be assumed to be fixed in place once they have nucleated. The evolution of the dilute phase follows equation (2.43).
@@ 4.4.2  Network permeation
If the elastic network is sufficiently stiff and/or the mesh size is sufficiently large, it may be energetically favourable for droplets to permeate, rather than exclude, the network. As droplets permeate the network, a further contribution to the free energy is induced, resulting from the differential surface energies of network–droplet and network–solvent interfaces. Following the work of Ronceray et al. [374], we capture this contribution by introducing a ‘wetting energy’ of the form
@@ 4.4.3  Complex properties of the mesh
Equation (4.10) highlights the dependence of the equilibrium droplet size on the elastic properties of the mesh, via the elastic pressure. Typically, droplets will be larger in softer materials. If the elastic mesh has long-range variations in its stiffness, we would thus expect larger droplets to form in softer regions of the mesh [134, 387]. Over long timescales, it is then energetically favourable for droplets in stiffer regions to dissolve in favour of droplet growth in softer regions, in order to equilibrate chemical potential across the system. This process has been termed elastic ripening [143], and is distinct from the classical Ostwald ripening effect. In particular, it has been shown that larger droplets can dissolve at the expense of smaller droplets in appropriate stiffness gradients, in contrast to Ostwald ripening [143].
@@ 5 Chemically active droplets
Cells can control the properties of biomolecules by chemical modifications, such as post-translational modifications in the case of proteins. Such modifications can affect physical interactions with other biomolecules and thus also phase separation. Many condensates in cells are regulated by chemical modifications [391, 224, 53, 223, 43], including centrosomes [392] and stress granules [393]. To develop a physical understanding of these regulatory mechanisms, we conceptualise biomolecular modifications as chemical transitions between two forms of the same molecule, differing in their interaction parameters encoded by $w_{i}$ and $\chi_{ij}$. In this section, we discuss the complex interplay between chemical reactions and phase separation. After introducing a general framework for describing (active) chemical reactions, we couple reactions to phase separation, identifying two distinct classes of chemically active droplets in the simplest case of a binary mixture.
@@ 5.1 Chemical reactions of many components
Figure 5.1: Qualitatively different classes of chemically active droplets. Reaction flux $s$ as a function of the volume fraction $\phi$ of droplet material for various qualitatively different functions $s(\phi)$. The flux diverges for extreme fractions (equation 5.6) and vanishes for either one fraction (panel A) or multiple fractions (panel B). Homogeneous states with fraction $\phi$ are stable with respect to reactions if $s^{\prime}(\phi)<0$, implying that only the outer two roots are stable in panel B. The relative positions of these roots in comparison to the equilibrium fractions $\phi^{(0)}_{\mathrm{in/out}}$ delimiting the blue binodal region determines the basal reaction rates $\Gamma_{\mathrm{in/out}}$, and thus the two distinct classes of externally maintained droplets (orange, $\Gamma_{\mathrm{in}}<0$, $\Gamma_{\mathrm{out}}>0$) and internally maintained droplets (green, $\Gamma_{\mathrm{in}}>0$, $\Gamma_{\mathrm{out}}<0$); all other cases lead to homogeneous systems (grey lines).
Reaction kinetics describe how the mixture composition $\Phi=(\phi_{1},\ldots,\phi_{N_{c}})$ changes over time. For a homogeneous mixture, there are no diffusive fluxes, implying $\partial_{t}\phi_{i}=s_{i}$, where $s_{i}=\sum_{\alpha=1}^{N_{\mathrm{r}}}\sigma^{(\alpha)}_{i}s^{(\alpha)}$ is the net production of component $i$. The net flux $s^{(\alpha)}$ of reaction $\alpha$ can be split into a forward and a backward direction, $s^{(\alpha)}=s^{(\alpha)}_{\rightarrow}-s^{(\alpha)}_{\leftarrow}$, which must obey detailed balance of the rates [23, 394],
@@
Detailed balance of the rates ensures that the reaction ceases ($s^{(\alpha)}=0$) when chemical equilibrium as described by equation (5.2) is reached, and also that the reaction proceeds in the forward direction when products are energetically favoured over reactants. A simple theory that obeys these constraints is transition state theory (appendix F and ref. [395, 396, 397, 398, 399, 87]),
$s^{(\alpha)}=k_{\alpha}(\Phi)\sinh\biggl{[}-\frac{\sum_{i}\sigma^{(\alpha)}_{i}\,\mu_{i}}{2k_{\mathrm{B}}T}\biggr{]}\;,$ (5.4)
@@
where the square bracket describes energetic constraints, whereas kinetic effects are captured by the positive rate coefficient $k_{\alpha}$, which can depend on the composition $\Phi$. The functional form of $k_{\alpha}$ can be constrained by requiring that $s^{(\alpha)}$ reduces to mass action kinetics for ideal solutions, implying that (appendix F)
$k_{\alpha}(\Phi)=\hat{k}_{\alpha}\prod_{i=1}^{N_{\mathrm{c}}}\phi_{i}^{\frac{1}{2}\bigl{(}\sigma^{(\alpha)}_{\rightarrow,i}+\sigma^{(\alpha)}_{\leftarrow,i}\bigr{)}}\;,$ (5.5)
with constant coefficient $\hat{k}_{\alpha}$. This form of $k_{\alpha}$ can also describe reaction pathways that include an enzyme $E$. Equation (5.5) then predicts that the reaction rate is scaled by the corresponding fraction $\phi_{E}$ if a single enzyme participates in the reaction (${\sigma^{(\alpha)}_{\rightarrow,E}=\sigma^{(\alpha)}_{\leftarrow,E}=1}$). Note, however, that biologically realistic reactions might exhibit more complex dependencies on the volume fractions.
@@
The reaction kinetics described by equation (5.4) generally drive the system towards equilibrium. However, biological systems are active and driven away from equilibrium by their metabolism [400]. A theoretical representation of metabolic processes involves the conversion of high-energy fuel molecules into low-energy waste molecules via chemical reactions. These reactions allow ordered structures within the cell to be maintained while increasing the total entropy of the cell and its environment. Note that although the energy difference between fuel and waste molecules is dissipated as heat, resulting temperature changes are typically negligible in cellular systems [98]. If fuel molecules are plentiful, they essentially provide an infinite particle reservoir with fixed chemical potential, also known as a chemostat. If the production of waste molecules does not significantly change their concentration in the environment, they can also be considered to be chemostatted. The energy difference between chemostatted species drives chemical reactions out of equilibrium [401, 402, 403]. This framework can also be used to describe replenishing systems, where waste molecules (e.g., ADP) are turned back into fuel molecules (ATP) by some mechanism (ATP synthase) whose energy input is not described. In both of these situations, the supplied energy is used locally to drive processes against their thermodynamic tendency, implying that the resulting system is an example of active matter [404].
@@
Chemical reactions also take place in heterogeneous systems, where the composition $\Phi$ varies in space and time. The general dynamics of such systems is discussed in section 2, where we already introduced the source term $s_{i}$ in equation (2.14) to account for chemical reactions. The combination of equation (2.14) with equation (5.4) for $s^{(\alpha)}$ implies that chemical reactions affect composition only locally, whereas spatial transport takes place via diffusion and advection.
@@ 5.2 Chemically active binary mixtures
To discuss the influence of chemical reactions on droplets, we now focus on the simplest case of a binary mixture comprising droplet material $D$ and precursor material $P$. The composition of the system is then quantified by the volume fraction $\phi(\boldsymbol{r},t)$ of droplet material, while precursor material occupies the fraction ${1-\phi}$. Thermodynamic properties of the mixture are captured by the exchange chemical potential $\bar{\mu}$; see section 2.1.3. If the two species can freely convert into each other, ${P\rightleftharpoons D}$, the system can in principle attain the minimum free energy (with ${\bar{\mu}=0}$; see figure 2.1a) at every point in space, implying a homogeneous mixture as the only possible equilibrium state.
More interesting dynamics arise when the system is driven out of equilibrium by introducing an additional chemical reaction, $P+F\rightleftharpoons D+W$, where the fuel $F$ and waste $W$ are chemostatted. For simplicity, we assume that fuel and waste distribute fast and are dilute, and so do not affect phase separation [405]. These components therefore just provide a chemical potential difference $\Delta\!\mu=\mu_{F}-\mu_{W}$, which drives an active reaction. The combined reaction flux $s$ then reads
@@
where $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are the respective reaction coefficients for the passive and active reactions. Both $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are positive and may depend on the composition $\phi$.
The chemical reaction flux given by equation (5.6) results in the creation of droplet material when the droplet fraction $\phi$ is very small; i.e., $s\rightarrow\infty$ in the limit $\phi\rightarrow 0$. This result emerges because the exchange chemical potential diverges ($\bar{\mu}\rightarrow-\infty$) as $\phi\rightarrow 0$ (see equation (2.17) and figure 2.1b). Conversely, $s\rightarrow-\infty$ as $\phi\rightarrow 1$, implying that droplet material is converted to precursor if droplet concentration is too high. As a result, the reaction flux $s$ can be approximated as an explicit function of the composition $\phi$. Recall from equation (2.23) that the dynamics of the binary mixture are then governed by
@@
where $\Lambda$ denotes the diffusive mobility. The limiting behavior of $s(\phi)$ for large and small volume fraction are constrained by the thermodynamic arguments discussed above, but the behaviour at intermediate volume fractions depends on the precise choice of parameters and in particular the composition dependence of $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$. Figure 5.1 shows various different functional forms for $s(\phi)$, which we categorise below based on the qualitative behaviour that they induce.